# Learning Unmasking Policies for Diffusion Language Models

### v2

@@ 3.2 Lightweight Confidence Policy Design
Having outlined the MDP above, we next describe our proposed implementation for the sampling policy $\pi_{\phi}$.
Confidence-based input
@@
Lightweight transformer
We introduce a small, learnable neural network $f_{\phi}$ that maps the vector of confidences ${\bm{c}}_{t}$ to a vector of unmasking scores (logits) ${\bm{b}}_{t}=f_{\phi}({\bm{c}}_{t},{\bm{m}}_{t},t)$ with ${\bm{b}}_{t}\in\mathbb{R}^{L}$. We additionally include a binary mask vector ${\bm{m}}_{t}:=(m_{t}^{1},\ldots,m_{t}^{L})$, where $m_{t}^{k}:=\mathbf{1}[k\in\mathcal{M}_{t}]$ indicates whether position $k$ is still masked, and inform the policy with the time index $t\in[T]$. In practice, $f_{\phi}$ is implemented as a lightweight transformer (see Section˜4 for more details on the exact policy architecture). Notably, its size is less than $0.01\%$ of the pretrained dLLMs used in our experiments, resulting in negligible computational overhead during sampling.
Bernoulli likelihood
We sample the unmasking actions according to $u_{t}^{k}\sim\text{Ber}\big(s_{t}^{k}\big)$ with $s_{t}^{k}:=\sigma(b_{t}^{k})$, where $\text{Ber}(\cdot)$ denotes the Bernoulli distribution and $\sigma(\cdot)$ is the sigmoid function. Conveniently, the policy likelihood $\pi_{\phi}({\bm{u}}_{t}):=\prod_{k=1}^{L}(s_{t}^{k})^{u_{t}^{k}}\cdot(1-s_{t}^{k})^{1-u_{t}^{k}}$ is readily available in closed form and does not require any additional approximations (unlike in the case of post-training dLLMs; zhao2025d1). We also considered an alternative formulation based on the Plackett-Luce model (luce1959individual; plackett1975analysis) which we call dynamic Plackett-Luce sampling (DPLS; see Appendix C), but since our ablations showed comparable performance (Section˜4.4), we favor the Bernoulli formulation in our main experiments for its simplicity.
At generation, we additionally propose to “temper” the Bernoulli parameters $s_{t}^{k}:=\sigma(b_{t}^{k}/\tau_{\pi})$ , where $\tau_{\pi}$ is the policy sampling temperature. This temperature controls the sharpness of the policy distribution and, as shown in Section˜4, we found that it can sometimes serve as a test-time knob for balancing performance (for example, $\tau_{\pi}<1$ forces the Bernoulli distribution to be more “decisive” by pushing the probability more towards either 0 or 1 depending on the sign of the logit $b_{t}^{k}$). Moreover, to ensure convergence when all sampled actions are zero (${\bm{u}}_{t}=\mathbf{0}$), we unmask the position with the highest Bernoulli parameter $s_{t}^{k}$ (similar to Fast-dLLM, see Section˜2.2). We apply this fallback only at generation time, as we found that forcing unmasking during training was prone to reward hacking, where the model would reliably obtain non-zero (but sub-optimal) reward simply by pushing all its predictions $b_{t}^{k}$ towards -$\infty$. By not enabling this behavior during training, the policy is forced to actively choose which tokens to unmask.
@@ 3.3 Learning $\pi_{\phi}$ with GRPO
where $\pi_{\phi}$ is the current policy being updated, and $\pi_{\phi_{\text{old}}}$ refers to an earlier version of the policy used to generate the RL rollouts. The likelihood ratio $\rho_{t}^{g}:=\frac{\pi_{\phi}({\bm{u}}_{t}^{g})}{\pi_{\phi_{\text{old}}}({\bm{u}}_{t}^{g})}$ serves as the importance sampling correction term, which together with clipping (via $\epsilon$) aims to stabilize the off-policy training. Since our policies are trained from scratch, we remove the KL regularization term in our GRPO objective. Finally, we note that GRPO can be equivalently interpreted as an off-policy variant of the standard REINFORCE algorithm (williams1992simple), where the group mean normalization functions as a control variate to reduce the variance of the policy gradient (mohamed2020monte).
@@
where $r({\bm{y}},{\bm{y}}_{t})$ is a task-specific correctness term (e.g., a binary reward indicating whether the generated mathematical answer is correct). To encourage faster sampling, we incorporate a computational penalty based on the number of steps, $T-\hat{T}$, with $\alpha\geq 0$ serving as a hyperparameter that controls the trade-off between accuracy and speed. While we experimented with an additive penalty of the form $r({\bm{y}}_{\hat{T}},{\bm{y}})-\alpha\big(\frac{T-\hat{T}}{T}\big)$ (graves2016adaptive) we found that this led to problematic reward hacking: when all samples in a group are incorrect, as is the case early on when training from scratch, faster samples may still receive a positive advantage, despite being wrong. Such additive reward derailed and destabilized RL training (cf. Section˜4.4). Our multiplicative reward can therefore prioritize first correctness, and then efficiency.
4 Experiments
We begin our experiments by verifying that the learned policies match the performance of confidence-based heuristics (Section˜4.1). We then highlight the potential of RL-based sampling to outperform heuristics in the more general setting without semi-AR decoding (Section˜4.2), and examine the transferability of our RL policies across models, datasets, and generation lengths (Section˜4.3). Finally, we explore different ways to instantiate the MDP and analyze their effects on downstream performance (Section˜4.4).
4.1 Learning Effective Sampling via RL
Experiment details.
We use LLaDA-8B-Instruct (llada2025) as the base dLLM (additional results on Dream-7B-Instruct (ye2025dream) provided in Appendix A.2). Note that while LLaDA is trained as an MDM from scratch, Dream is initialized from an AR model (Qwen 2.5; qwen2025qwen25technicalreport). We parametrize the policy network $f_{\phi}$ as a shallow (single-layer) transformer incorporating adaptive layer normalization for conditioning; hyperparameter details for the architecture are provided in Appendix˜B. We train five different policies, corresponding to $\alpha\in\{10,3,1,0.3,0\}$. Each policy is trained semi-autoregressively at $BL=32$ on a single epoch of mixture data, sampled proportionally from the training sets of GSM8k (cobbe2021gsm8k) and MATH (hendrycksmath2021), resulting in roughly 15,000 training samples. Figure˜2 shows the training dynamics for these runs; as expected, higher $\alpha$ generally gives rise to faster policies.
We then compare the resulting policies on the test sets of both GSM8k and MATH to the confidence-based heuristics introduced in Section˜2.2, averaging over two training seeds and three test-time seeds. To obtain Pareto frontiers for the heuristic methods, we use $K\in\{8,16,32,64,128,256\}$ for the random baseline and high-confidence unmasking, while for Fast-dLLM we use $\lambda\in\{0.1,0.2,\ldots,1.0\}$ throughout. For all methods, we use the standard greedy decoding setting ($\tau=0$) when generating test answers. To measure sampling speed, we report the mean number of function evaluations (NFEs) of the underlying dLLM.
Results with (Short) 32 Block Length, 3(a) and 3(c)
Generally, we observe that the performance of the learned policies ( ) exceeds those of both the random baseline ( ) as well as the high-confidence sampling ( ) strategy, while matching that of Fast-dLLM ( ). The fact that our learned policies do not surpass Fast-dLLM in the mid-to-high NFEs range may suggest that this heuristic is near-optimal under the semi-AR sampling regime, where spatial correlation is enforced during the decoding process.
We note that the effect of $\alpha$ in training (cf. Figure˜2) carries over to the generation on test set: policies trained with higher values of $\alpha$ exhibit faster, yet less accurate, sampling. Of particular note is $\alpha=10.0$, which exhibited greater training instability with only 1 out of 2 training runs converging. When the seed that does take off is evaluated, we find that it yields a very fast policy that outperforms Fast-dLLM in the low-NFE ($\sim 10$) regime on both datasets. This highlights the potential of RL policies when optimizing for maximal efficiency. However, RL policies appear to exhibit less controllability compared to a heuristic like Fast-dLLM, as varying $\alpha$ results in a less smooth traversal of the Pareto frontier than varying the confidence threshold $\lambda$. For example, the behavior of the $\alpha=3.0$ policy is nearly identical to that of $\alpha=1.0$, despite the reward function incorporating a computational penalty that scales exponentially with $\alpha$. Furthermore, we find that this cannot easily be remedied by using a tighter grid for $\alpha$. In Appendix A.3 we let $\alpha$ range in $\{10.0,9.0,...1.0,0.3,0.0\}$, but observe that using $\alpha\geq 4.0$ consistently results in converging to a policy either roughly equivalent to that of $\alpha=3.0$ or to that of $\alpha=10.0$, with no value of $\alpha$ successfully yielding a policy which interpolates between the two extremes. Besides $\alpha$, we observe that one can make small changes to the trade-off between compute and performance by varying the policy temperature $\tau_{\pi}$ at test time; we detail the impact of this parameter in Figure 9.
4.2 Beyond Semi-Autoregressive Decoding
As mentioned in Section˜2.2, heuristic sampling methods often rely on semi-AR generation to achieve good performance. While not typically thought of as problematic in textual domains where strong AR dependencies exist, semi-AR approaches can only partially fulfill the promise of fully parallel generation. Having verified that RL can help learn sampling policies with short block-length, we consider in section longer blocks.
@@ 4.2 Beyond Semi-Autoregressive Decoding
Figure 3: Results for LLaDA in semi-AR (Figure˜3(a) & Figure˜3(c)) and full-diffusion (Figure˜3(b) & Figure˜3(d)) generation regimes. Results for Dream-7B are provided in Figure 7. For our policies we vary $\alpha\in\{10,3,1,0.3,0\}$ and use $\tau_{\pi}=0.5$ for $BL=32$ and $\tau_{\pi}=1$ for $BL=256$. Expert steering (ES) described in detail in Appendix D.
@@
We train policies without relying on semi-AR generation, targeting the full-length generation task ($BL=L=256$) at both training and test time. We keep all other implementation details identical to those in the previous section (Section˜4.1), and evaluate on the test sets of GSM8k and MATH.
Results with (Long) 256 Block Length, 3(b) & 3(d)
@@
Although all sampling strategies experience a performance drop relative to the semi-AR setting (cf. 3(a) & 3(c)), we find that the policies produced by RL ( ) exhibit the smallest decline and consequently achieve the best overall performance. Furthermore, the results suggest that this methodology is able to produce policies which achieve solid performance even in the low-NFE regime. In particular, the learned policies obtain $\sim 50\%$ accuracy at $\sim 12$ NFEs on GSM8k (compared to $\leq 30\%$ for the heuristic methods regardless of semi-AR use), highlighting the potential of non-semi-AR sampling for achieving maximal efficiency gains. Note however that, in theory, a RL policy trained with $BL=L$ could learn to emulate semi-AR sampling on its own. Thus, the fact that policies trained with $BL=L$ underperform those trained with $BL=32$ in the mid-high NFE range (comparing, for example, Figure˜3(a) to Figure˜3(b)) suggests that the policies remain far from optimal. One hypothesis for why this is the case is that our training procedure is not encouraging sufficient exploration, instead converging to locally optimal policies. To address this, we explore encouraging further exploration by using samples generated via semi-AR sampling from Fast-dLLM. We refer to this approach as expert steering, and describe it in more detail in Appendix D. As shown in Figure˜3, with expert steering ( ) RL is able to discover policies that close the performance gap to the best accuracy achieved in the semi-AR setting at mid-to-high NFEs (e.g., $\sim 80\%$ on GSM8k and $\sim 35\%$ on MATH), while mostly retaining their strong performance in the low-NFE regime. However, we do find that expert steering introduces significant instability during training, further reducing the controllability through $\alpha$, with multiple values of $\alpha$ collapsing to near-identical policies. We leave further investigations into stabilizing expert steering training for future work.

### v3

@@ 3.2 Lightweight Confidence Policy Design
Figure 2: Diagram showing how our policies are used on top of a pretrained dLLM to unmask tokens and generate text.
Having outlined the MDP above, we next describe our proposed implementation for the sampling policy $\pi_{\phi}$ (see Figure˜2 for policy sampling diagram and Appendix A for sampling algorithm).
Confidence-based input.
@@
Lightweight transformer.
We introduce a small, learnable neural network $f_{\phi}$ that maps the vector of confidences ${\bm{c}}_{t}$ to a vector of unmasking scores (logits) ${\bm{b}}_{t}=f_{\phi}({\bm{c}}_{t},{\bm{m}}_{t},t)$ with ${\bm{b}}_{t}\in\mathbb{R}^{L}$. We additionally include a binary mask vector ${\bm{m}}_{t}:=(m_{t}^{1},\ldots,m_{t}^{L})$, where $m_{t}^{k}:=\mathbf{1}[k\in\mathcal{M}_{t}]$ indicates whether position $k$ is still masked, and inform the policy with the time index $t\in[T]$. In practice, $f_{\phi}$ is implemented as a lightweight transformer (see Section˜4 for more details on the exact policy architecture). Notably, its size is less than $0.01\%$ of the pretrained dLLMs used in our experiments, resulting in negligible computational overhead during sampling (see Figure˜10).
Bernoulli likelihood.
We sample the unmasking actions via $u_{t}^{k}\sim\text{Ber}\big(s_{t}^{k}\big)$ with $s_{t}^{k}:=\sigma(b_{t}^{k})$, where $\text{Ber}(\cdot)$ denotes the Bernoulli distribution and $\sigma(\cdot)$ is the sigmoid function. Conveniently, the policy likelihood $\pi_{\phi}({\bm{u}}_{t}):=\prod_{k=1}^{L}(s_{t}^{k})^{u_{t}^{k}}\cdot(1-s_{t}^{k})^{1-u_{t}^{k}}$ is readily available in closed form and does not require any additional approximations (unlike in the case of post-training dLLMs; zhao2025d1). We also considered an alternative formulation based on the Plackett-Luce model (luce1959individual; plackett1975analysis) which we call dynamic Plackett-Luce sampling (DPLS; see Appendix E), but since our ablations showed comparable performance (Section˜4.4), we favor the Bernoulli formulation in our experiments for its simplicity.
At generation, we additionally propose to “temper” the Bernoulli probabilities as $s_{t}^{k}:=\sigma(b_{t}^{k}/\tau_{\pi})$. The temperature $\tau_{\pi}$ thus controls the sharpness of the policy distribution; we found that this can sometimes serve as a useful test-time knob for trading off accuracy versus efficiency by making the policy more or less “decisive” (Section˜4).
To ensure convergence when all sampled actions are zero (${\bm{u}}_{t}=\mathbf{0}$), we unmask the position with the highest Bernoulli parameter $s_{t}^{k}$ (similar to Fast-dLLM). We apply this fallback only at test time, as we found that forcing unmasking during training was prone to reward hacking. Instead, if not all tokens are unmasked within $T$ steps during training, we simply terminate generation and evaluate the answer with some tokens left unmasked. Since such samples tend to get low or no reward, we find that the policies quickly learn not to leave any tokens unmasked without explicit instruction.
@@ 3.3 Learning $\pi_{\phi}$ with GRPO
where $\pi_{\phi}$ is the current policy being updated, and $\pi_{\phi_{\text{old}}}$ refers to an earlier version of the policy used to generate the RL rollouts. The likelihood ratio $\rho_{t}^{g}:=\frac{\pi_{\phi}({\bm{u}}_{t}^{g})}{\pi_{\phi_{\text{old}}}({\bm{u}}_{t}^{g})}$ serves as the importance sampling correction term, which together with clipping (via $\epsilon$) aims to stabilize the off-policy training. Note that we exclude already-unmasked positions from the policy likelihood computation. Finally, since our policies are trained from scratch, we remove the KL regularization term in our GRPO objective.
@@
where $r({\bm{y}},{\bm{y}}_{t})$ is a task-specific correctness term (e.g., a binary reward indicating whether the generated mathematical answer is correct). To encourage faster sampling, we incorporate a computational penalty based on the number of steps, $T-\hat{T}$, with $\alpha\geq 0$ serving as a hyperparameter that controls the trade-off between accuracy and speed. While we experimented with an additive penalty of the form $r({\bm{y}},{\bm{y}}_{\hat{T}})-\alpha\big(\frac{T-\hat{T}}{T}\big)$ (graves2016adaptive) we found that this led to problematic reward hacking: when all samples in a group are incorrect, as is the case early on when training from scratch, faster samples may still receive a positive advantage, despite being wrong (cf. Section˜4.4).
@@
Figure 4: Results for LLaDA in semi-AR (Figure˜4(a) & Figure˜4(c)) and full-diffusion (Figure˜4(b) & Figure˜4(d)) generation regimes ($L=256$). Results for Dream-7B are provided in Figure 11. For our policies we vary $\alpha\in\{10,3,1,0.3,0\}$ and use $\tau_{\pi}=0.5$ for $BL=32$ and $\tau_{\pi}=1$ for $BL=256$. Expert steering (ES) described in detail in Appendix F. Wall-clock time plots shown in Figure˜10.
4 Experiments
We begin by verifying that the learned policies match the performance of confidence-based heuristics (Section˜4.1). We then highlight the potential of RL sampling to outperform heuristics in the full-diffusion setting (Section˜4.2), and examine the transferability of RL policies across models, datasets, and generation lengths (Section˜4.3). Next, we explore different ways to instantiate the MDP and analyze their effects on performance (Section˜4.4). Finally, we study the unmasking behavior of RL policies, highlighting their qualitative differences to heuristics (details in Appendix C).
4.1 Learning Effective Sampling via RL
@@
We use LLaDA-8B-Instruct (llada2025) as the base dLLM (additional results on Dream-7B-Instruct (ye2025dream) provided in Appendix B.3). We parametrize the policy network $f_{\phi}$ as a shallow (single-layer) transformer incorporating adaptive layer normalization for conditioning; hyperparameter details for the architecture are provided in Appendix H and an architecture diagram can be found in Appendix I. We train five different policies, corresponding to $\alpha\in\{10,3,1,0.3,0\}$. Each policy is trained semi-autoregressively at $BL=32$ on a single epoch of mixture data, sampled proportionally from the training sets of GSM8k (cobbe2021gsm8k) and MATH (hendrycksmath2021), resulting in roughly 15,000 training samples. Figure˜3 shows the training dynamics; as expected, higher $\alpha$ generally gives rise to faster policies.
We then compare the resulting policies on the test sets of both GSM8k and MATH to the confidence-based heuristics introduced in Section˜2.2, picking one out of two training seeds based on the final training loss and averaging over three test-time seeds. To obtain Pareto frontiers for the heuristic methods, we use $K\in\{8,16,32,64,128,256\}$ for the random baseline and high-confidence unmasking, while for Fast-dLLM we use $\lambda\in\{0.1,0.2,\ldots,1.0\}$ throughout. For all methods, we use the standard greedy decoding setting ($\tau=0$) when generating test answers.
Results for (Short) $BL=32$, 4(a) and 4(c).
Generally, we observe that the performance of the learned policies ( ) exceeds those of both the random baseline ( ) as well as the high-confidence sampling ( ), while matching Fast-dLLM ( ). That our learned policies do not surpass Fast-dLLM in the mid-to-high NFEs range may suggest that this heuristic is near-optimal under semi-AR generation. Interestingly, despite very similar quantitative performance, we find that the learned policies exhibit qualitatively somewhat different unmasking behavior compared to Fast-dLLM (see Appendix C). Concretely, the policy unmasks adjacent tokens much less frequently (Figure˜20) and allocates compute more uniformly across blocks (Figure˜5(a)).
(a) BL32-slow ( )
(b) BL32-fast ( )
Figure 5: Average number of sampling steps (NFEs) per block for LLaDA on GSM8k with semi-AR generation (same setting as in Figure˜4(a)). Left we show NFEs per block for ‘slow’ variants of Fast-dLLM and policy sampling ($\sim 75$ NFEs), while in the right plot we show ‘fast’ variants ($\sim 10$ NFEs). Fast-dLLM exhibits a pattern of allocating more compute to earlier blocks. Our policy sampling, by contrast, distributes compute more uniformly across blocks, except in the fast ($\alpha=10$) policy, where most compute is expended in the final block while generating numerical answers. To see where evaluated policies here are located on the pareto frontier, see Figure˜17(a). More details on qualitative differences between Fast-dLLM and policy sampling are provided in Appendix C.
We note that the effect of $\alpha$ in training (cf. Figure˜3) carries over to the test set: policies trained with higher values of $\alpha$ exhibit faster, yet less accurate, sampling. Of particular note is $\alpha=10.0$, which exhibited greater training instability due to the very sharp slope of the reward. When successfully trained, this yields a very fast policy that outperforms Fast-dLLM in the low-NFE ($\sim 10$) regime on both datasets. Qualitatively, we find evidence that this improved performance is likely due to the policy learning to slow down in the final block when generating a numerical answer, as shown in Figure˜5(b) (see also Appendix C for more details). This highlights the potential of RL-based policies when optimizing for maximal efficiency.
However, RL policies appear to exhibit less controllability, as varying $\alpha$ results in a less smooth traversal of the Pareto frontier than varying the confidence threshold $\lambda$. For example, the behavior of the $\alpha=3.0$ policy is nearly identical to that of $\alpha=1.0$, despite the reward incorporating a computational penalty that scales exponentially with $\alpha$. Furthermore, we find that this cannot easily be remedied by using a tighter $\alpha$ grid. In Appendix B.4 we let $\alpha$ range in $\{10.0,9.0,...1.0,0.3,0.0\}$, but observe that using $\alpha\geq 4.0$ consistently results in converging to a policy either roughly equivalent to that of $\alpha=3.0$ or to that of $\alpha=10.0$, with no value of $\alpha$ successfully yielding a policy which interpolates between the two extremes. Besides $\alpha$, we observe that one can make small changes to the trade-off between compute and performance by varying the policy temperature $\tau_{\pi}$ at test time; we detail the impact of this parameter in Figure 14.
4.2 Beyond Semi-Autoregressive Decoding
As mentioned in Section˜2.2, heuristic samplers often rely on semi-AR generation to achieve good performance. While not typically thought of as problematic in textual domains where strong AR dependencies exist, semi-AR approaches can only partially fulfill the promise of fully parallel generation by restricting the unmasking to the current block. Hence, we train here policies without semi-AR generation, targeting the full-diffusion task ($BL=L=256$) at both training and test time. We keep all other implementation details identical to those in the previous section.
Results for (Long) $BL=256$, 4(b) & 4(d).
@@ 4.2 Beyond Semi-Autoregressive Decoding
Although all sampling strategies experience a performance drop relative to the semi-AR setting (cf. 4(a) & 4(c)), we find that the policies produced by RL ( ) exhibit the smallest decline and consequently achieve the best overall performance. Furthermore, the results suggest that this methodology is able to produce policies which achieve solid performance even in the low-NFE regime. In particular, the learned policies obtain $\sim 50\%$ accuracy at $\sim 12$ NFEs on GSM8k (compared to $\leq 30\%$ for the heuristic methods regardless of semi-AR use), highlighting the potential of non-semi-AR sampling for achieving maximal efficiency gains. Note however that, in theory, an RL policy trained with $BL=L$ could learn to emulate semi-AR sampling on its own. Thus, the fact that policies trained with $BL=L$ underperform those trained with $BL=32$ in the mid-high NFE range (comparing, for example, Figure˜4(a) to Figure˜4(b)) suggests that the policies remain far from optimal.
One hypothesis for why this is the case is that our training procedure is not encouraging sufficient exploration, instead converging to locally optimal policies. To address this, we encourage further exploration by using samples generated via semi-AR sampling from Fast-dLLM. We refer to this approach as expert steering, and describe it in more detail in Appendix F. As shown in Figure˜4, with expert steering ( ) RL is able to discover policies that almost close the performance gap to the best accuracy achieved in the semi-AR setting at mid-to-high NFEs (e.g., $\sim 80\%$ on GSM8k and $\sim 35\%$ on MATH), while mostly retaining their strong performance in the low-NFE regime. However, we do find that expert steering introduces significant instability during training, further reducing the controllability through $\alpha$, with multiple values of $\alpha$ collapsing to near-identical policies. We leave further investigations into stabilizing expert steering training for future work. Finally, we observe that policy sampling and Fast-dLLM induce markedly different unmasking orders in the full-diffusion
Figure 6: Mean unmasking time for each token position ($L=256$), averaged over $N=100$ samples, for LLaDA on GSM8k under full-diffusion generation (same setting as in Figure˜4(b)). For visualization purposes, time is shown in reverse. Fast-dLLM ( ) exhibits somewhat reverse (i.e., right-to-left) generation due to LLaDA’s corrupted confidence on padding tokens (see Appendix C). Encouragingly, the expert steering policy ( ) learns to overcome this issue and recovers left-to-right generation (observe how tokens at earlier positions are generated first on average). The locations of the evaluated policies on the Pareto frontier are shown in Figure˜17(b) (see ).
setting (see Figure˜6 and Appendix C). While Fast-dLLM exhibits reverse (right-to-left) unmasking, the policy with expert steering learns to overcome this behavior and unmasks earlier positions first on average (left-to-right), which likely explains its superior performance.